# Mixup Model Merge: Enhancing Model Merging Performance through Randomized Linear Interpolation

### v1

@@ 4.1 Experimental Setup
The Benchmarks for evaluating Out-of-Distribution and Adversarial Robustness
To assess OOD robustness, we evaluate math & code, LM & math, and LM & code models using instruction following (LiveBench-Instruction), coding (LiveBench-Coding), and language comprehension (LiveBench-TypoFixing) category in LiveBench (White et al., 2024), respectively. More details on OOD benchmarks are given in Appendix A.3.
We utilize the Adversarial Prompt Attacks module in PromptBench (Zhu et al., 2024) to assess the robustness of LLMs against adversarial prompts. Specifically, we employ three attack methods: DeepWordBug (character-level) (Gao et al., 2018), BERTAttack (word-level) (Li et al., 2020), and StressTest (sentence-level) (Naik et al., 2018). The evaluation is conducted on two datasets supported by PromptBench: SST2 (sentiment analysis) (Socher et al., 2013) and CoLA (grammatical correctness) (Warstadt, 2019). For more details on PromptBench and attack methods, please refer to Appendix A.4.
@@
Unless otherwise specified, the details of the model merging experiments are consistent with Yu et al. (2024). The hyperparameter $\alpha$ for M3 is chosen from the range $[0.2,0.4,0.5,1,2,3,5]$. For a detailed description of the hyperparameter settings in model merging methods, please refer to Appendix A.2. Additionally, all experiments are conducted on NVIDIA GeForce RTX 4090 GPUs.
@@ 4.2 Merging Task-Specific Fine-Tuned LLMs
Merging Methods Models Use Model Mixup Use DARE Instruction Following Mathematical Reasoning Code Generating
AlpacaEval GSM8K MATH HumanEval MBPP
/ LM No No 45.14 2.20 0.04 36.59 34.00
Math No No / 64.22 14.02 / /
Code No No / / / 23.78 27.60
Average Merging LM & Math No No 45.28 66.34 13.40 / /
Yes No 44.40 66.26 13.80 / /
No Yes 44.22 66.57 12.96 / /
Yes Yes 43.53 66.57 14.12 / /
LM & Code No No 36.60 / / 29.88 32.00
Yes No 43.91 / / 37.20 34.40
No Yes 38.81 / / 31.71 32.40
Yes Yes 40.31 / / 36.59 37.00
Math & Code No No / 56.17 10.28 8.53 8.20
Yes No / 63.61 14.02 8.54 19.20
No Yes / 56.18 10.28 6.10 7.80
Yes Yes / 64.97 13.54 9.76 21.20
Task Arithmetic LM & Math No No 45.78 66.34 13.40 / /
Yes No 41.65 66.34 13.74 / /
No Yes 49.00 66.64 13.02 / /
Yes Yes 44.90 67.32 13.74 / /
LM & Code No No 44.64 / / 32.93 33.60
Yes No 46.64 / / 35.37 33.80
No Yes 41.47 / / 35.98 33.00
Yes Yes 45.20 / / 35.98 35.20
Math & Code No No / 64.67 13.98 8.54 8.60
Yes No / 63.53 13.94 7.93 19.00
No Yes / 65.05 13.96 10.37 9.80
Yes Yes / 65.13 14.32 8.54 18.00
TIES- Merging LM & Math No No 38.63 14.56 2.12 / /
Yes No 38.73 18.57 2.48 / /
No Yes 37.92 18.04 2.34 / /
Yes Yes 39.93 19.26 2.82 / /
LM & Code No No 41.85 / / 0.0 0.0
Yes No 44.96 / / 25.61 30.80
No Yes 43.13 / / 0.0 0.0
Yes Yes 45.65 / / 26.83 33.20
Math & Code No No / 64.67 13.68 9.15 22.60
Yes No / 64.75 14.16 9.76 21.4
No Yes / 64.82 13.88 10.37 23.60
Yes Yes / 64.75 14.78 9.15 19.60
Table 1: Performance of merging task-specific LLMs WizardLM-13B (LM), WizardMath-13B (Math), and llama-2-13b-codealpaca (Code) on all the datasets. The best and second-best results are marked in bold and underlined fonts.
We integrate M3 into three prominent model merging techniques: Average Merging, Task Arithmetic, and TIES-Merging. The performance of merging task-specific fine-tuned LLMs is presented in Table 1.
From Table 1, we obtain the following observations: 1) M3 generally enhances Average Merging, Task Arithmetic, and TIES-Merging when merging fine-tuned LLMs. For example, the improvements achieved by Average Merging with M3 for Math & Code are 7.43% on GSM8K, 3.74% on Math, and 11.0% on MBPP. For LM & Code, Average Merging with M3 shows improvements of 7.31% on AlpacaEval, 7.32% on HumanEval, and 2.4% on MBPP. Task Arithmetic with M3 results in improvements of 2.0% on AlpacaEval and 2.44% on HumanEval for LM & Code, and 10.4% on MBPP for Math & Code. TIES-Merging with M3 achieves an improvement of 4.01% for LM & Math on GSM8K. For LM & Code, TIES-Merging with M3 shows significant improvements of 3.11% on AlpacaEval, 25.61% on HumanEval, and 30.8% on MBPP. 2) Compared to Task Arithmetic, Average Merging and TIES-Merging tend to benefit more from M3. This is because both Average Merging and TIES-Merging use a fixed merging ratio of 1/2, whereas Task Arithmetic allows the merging ratio to vary within the range $[0.5,1.0]$. Consequently, the randomness introduced by M3 in the merging ratio has a more pronounced impact on Average Merging. This further highlights the critical role of an effective merging ratio in determining the performance of the merged model. 3) Yu et al. (2024) has indicated that the suboptimal results of merging WizardMath-13B with llama-2-13b-code-alpaca are due to llama-2-13b-code-alpaca not being well fine-tuned for code generation. In this context, the proposed M3 approach improves the pass@1 score on MBPP by 10.4% for the merged model of WizardMath-13B and llama-2-13b-code-alpaca. The improvement demonstrates that when one of the fine-tuned models to be merged is not well fine-tuned for the specific task, M3 can effectively unlock the potential of both models, maximizing the performance of the merged model. The M3 approach helps mitigate the impact of suboptimal fine-tuning on model merging performance.
@@ 4.3 Model Robustness
To ensure that the evaluation datasets are as representative as possible of OOD data, we select datasets with sufficiently recent release dates and ensure they cover domains that fine-tuned LLMs have not been specifically trained on. Consequently, Math & Code is evaluated on LiveBench-Instruction, LM & Math on LiveBench-Coding, and LM & Code on LiveBench-TypoFixing. The performance of the merged LLMs is shown in Figure 4. As illustrated in Figure 4, M3 consistently enhances the performance of merged models—Math & Code, LM & Math, and LM & Code—on OOD datasets. Specifically, when Task Arithmetic is combined with M3, the Math & Code model demonstrates a 1.9% improvement in win rate on LiveBench-Instruction, the LM & Math model achieves a 1.6% increase in pass@1 on LiveBench-Coding, and the LM & Code model shows a significant 6% boost in accuracy on LiveBench-TypoFixing. Similarly, when Average Merging is combined with M3, the Math & Code model attains a 1.5% improvement in win rate on LiveBench-Instruction, the LM & Math model achieves a 0.7% increase in pass@1 on LiveBench-Coding, and the LM & Code model exhibits a 4% enhancement in accuracy on LiveBench-TypoFixing. Finally, when TIES-Merging is applied alongside M3, the Math & Code model achieves a 1.1% improvement in win rate on LiveBench-Instruction, the LM & Math model records a 0.6% increase in pass@1 on LiveBench-Coding, and the LM & Code model demonstrates a remarkable 14% improvement in accuracy on LiveBench-TypoFixing. These results underscore the robustness and versatility of M3 in enhancing model performance across diverse merging strategies and OOD tasks.
@@
We employ three Prompt Attack Methods supported by the promptbench codebase (DeepWordBug, BERTAttack, and StressTest) (Zhu et al., 2024) to evaluate the adversarial robustness of three merged models (Math & Code, LM & Math, and LM & Code) obtained through the task arithmetic method. To balance experimental effectiveness with computational efficiency, we randomly selected the positions of the three attacked words in the prompts when executing the DeepWordBug and BERTAttack attacks. Adversarial robustness is assessed using the Performance Drop Rate (PDR) (Zhu et al., 2023), where a lower PDR indicates stronger robustness. Further details on PDR can be found in Appendix D. The performance of the merged LLMs is shown in Table 2.
@@
4.4 Mixup Model Merge with DARE
DARE is a model sparsification method proposed by (Yu et al., 2024), with a more detailed introduction provided in Appendix E. We combine M3 and DARE with three model merging techniques, including Average Merging, Task Arithmetic, and TIES-Merging, to compare the effects of M3 and DARE individually and explore their combined impact. The experimental results are presented in Table 1. Additionally, in the DARE method, the drop rate hyperparameter is set to 0.2.
From Table 1, we conclude that: 1) In most cases, M3 outperforms DARE, with a particularly significant advantage on certain datasets. For instance, the Math & Code model achieves a pass@1 score of 9.8% on the MBPP dataset when combined with DARE, while this score increases to 19% when combined with M3. This demonstrates that M3 unlocks new potential in model merging by randomly generating merging ratios, leading to performance improvements that surpass those achieved by DARE. 2) Combining DARE and M3 generally results in better model merging performance. For example, the LM & Math and LM & Code models, enhanced by TIES-Merging with M3 and DARE, achieve the best performance on the test datasets. While only incorporating TIES-Merging with M3 to these models, the enhanced models achieve the second best performance. This suggests that M3 and DARE can complement each other. Moreover, in some cases, M3 alone can deliver the best results, while DARE alone only achieves the best performance in very few cases, further demonstrating the superiority of M3.
@@ 5 Conclusion
Inspired by the mixup method and the Vicinal Risk Minimization (VRM) principle, we propose Mixup Model Merge (M3), a novel approach for merging fine-tuned LLMs by introducing randomness into the parameters linear interpolation process. Unlike traditional methods such as average merging and task arithmetic, M3 leverages a Beta distribution to dynamically adjust the merging ratio, enabling more flexible exploration of the parameter space. Experimental results demonstrate that M3 not only significantly enhances the performance of the merged model across various tasks but also improves its OOD and adversarial robustness. Furthermore, when combined with sparsification techniques such as DARE, our approach achieves even more favorable model merging outcomes. In summary, M3 is a simple yet powerful technique that requires minimal computational resources. By merely adjusting the merging ratio, it produces a merged model with enhanced task-specific capabilities and robustness. This exciting discovery paves the way for further research into optimizing merging ratio selection in model merging processes.
@@ 6 Limitations
There are several limitations of the M3 method: While it performs well for merging two models, (1) its scalability when merging a larger number of models, especially those with significant differences, remains uncertain. Additionally, (2) due to the inherent randomness in the merging process, multiple attempts may be required to achieve a merged model that meets expectations. This unpredictability can lead to increased computational costs, particularly in large-scale applications, resulting in a significant rise in resource consumption. Finally, (3) Our method may also be extended to a wider range of applications, such as merging fine-tuned models with Reinforcement Learning with Human Feedback (RLHF) (Ouyang et al., 2022) models to reduce the alignment tax (FINE-TUNING, ).

### v2

@@ 4.1 Experimental Setup
Benchmarks for evaluating Out-of-Distribution and Adversarial Robustness
To evaluate OOD robustness, we evaluate merged models including math & code, LM & math, and LM & code on instruction following (LiveBench-Instruction), coding (LiveBench-Coding), and language comprehension (LiveBench-TypoFixing) category in LiveBench (White et al., 2024), respectively. More details on OOD benchmarks are given in Appendix A.3.
We utilize the Adversarial Prompt Attacks module in PromptBench (Zhu et al., 2024) to evaluate the robustness of LLMs against adversarial prompts. Specifically, we employ three attack methods: DeepWordBug (character-level) (Gao et al., 2018), BERTAttack (word-level) (Li et al., 2020), and StressTest (sentence-level) (Naik et al., 2018). The evaluation is conducted on two datasets supported by PromptBench: SST2 (sentiment analysis) (Socher et al., 2013) and CoLA (grammatical correctness) (Warstadt, 2019). For more details on PromptBench and attack methods, please refer to Appendix A.4.
@@
In each distinct merging experiment—where “distinct” refers to variations in either the task-specific LLMs or the model merging methods—we perform a full sweep over the sole hyperparameter of M3, $\alpha$, using the set $\{0.2,0.4,0.5,1,2,3,5\}$. For each $\alpha$ value, we sample $\lambda_{m}$ once from the corresponding Beta distribution, resulting in a total of seven samples in each merging experiment. Seven samples are not the minimum required to observe significant performance gains. In most cases, a smaller number of samples is sufficient to identify a performant interpolation coefficient. However, to enable a more comprehensive exploration, we fix the number of samples to seven—each corresponding to a representative shape of the Beta distribution. While the stochastic nature of our method means there may be cases where none of the seven samples yields a significant improvement, such cases are not often observed in our experiments and thus not considered. Given the strong performance observed with just seven samples, and to conserve computational resources, we refrain from further sampling.
Unless otherwise specified, the remaining details of the model merging experiments follow Yu et al. (2024). For a detailed description of the hyperparameter settings used in the representative existing model merging methods, please refer to Appendix A.2. All experiments are conducted on NVIDIA GeForce RTX 4090 GPUs.
@@ 4.2 Merging Task-Specific Fine-Tuned LLMs
Merging Methods Models Use M3 (Ours) Sampling Details Instruction Following Mathematical Reasoning Code Generating Avg.
$\lambda_{m}$ $\alpha$ AlpacaEval GSM8K MATH HumanEval MBPP
- LM - - - 45.14 2.20 0.04 36.59 34.00 -
Math - - - - 64.22 14.02 - - -
Code - - - - - - 23.78 27.60 -
Average Merging LM & Math No - - 45.28 66.34 13.40 - - 41.67
Yes 0.35 2 44.40 66.26 13.80 - - 41.48
LM & Code No - - 36.60 - - 29.88 32.00 32.82
Yes 0.84 2 43.91$\uparrow$ - - 37.20$\uparrow$ 34.40 38.50$\uparrow$
Math & Code No - - - 56.17 10.28 8.53 8.20 20.80
Yes 0.92 0.4 - 63.61$\uparrow$ 14.02$\uparrow$ 8.54 19.20$\uparrow$ 26.34$\uparrow$
Task Arithmetic LM & Math No - - 45.78 66.34 13.40 - - 41.84
Yes 0.34 2 41.65 66.34 13.74 - - 40.58
LM & Code No - - 44.64 - - 32.93 33.60 37.06
Yes 0.99 0.4 46.64 - - 35.37 33.80 38.60
Math & Code No - - - 64.67 13.98 8.54 8.60 23.95
Yes 0.98 0.5 - 63.53 13.94 7.93 19.00$\uparrow$ 26.1
TIES- Merging LM & Math No - - 38.63 14.56 2.12 - - 18.44
Yes 0.62 1 38.73 18.57$\uparrow$ 2.48 - - 19.92
LM & Code No - - 41.85 - - 0.0 0.0 13.95
Yes 0.84 0.5 44.96 - - 25.61$\uparrow$ 30.80$\uparrow$ 33.79$\uparrow$
Math & Code No - - - 64.67 13.68 9.15 22.60 27.57
Yes 0.63 3 - 64.75 14.16 9.76 21.4 27.52
DARE LM & Math No - - 49.00 66.64 13.2 - - 42.95
Yes 0.38 0.4 44.90 67.32 13.74 - - 41.98
LM & Code No - - 41.47 - - 35.98 33.00 36.82
Yes 0.99 0.5 45.20$\uparrow$ - - 35.98 35.20 38.79
Math & Code No - - - 65.05 10.37 9.15 9.80 23.59
Yes 0.98 0.5 - 65.13 14.32$\uparrow$ 8.54 18.00$\uparrow$ 26.50$\uparrow$
Table 2: Comparison of Merged Model Performance Before and After Applying M3. We use three task-specific LLMs: WizardLM-13B (LM), WizardMath-13B (Math), and llama-2-13b-codealpaca (Code). Bold indicates the best results for each dataset, underline highlights improvements with M3, and $\uparrow$ denotes significant gains after introducing M3.
Our proposed method, M3, is a plug-and-play approach. To evaluate its effectiveness in improving merged model performance, we integrate M3 into several representative existing model merging methods (Average Merging, Task Arithmetic, TIES-Merging, DARE) and compare the performance of the merged models before and after applying M3. The results is presented in Table 2.
Performance Gains of M3 Across Merging Methods and Tasks
As shown in Table 2, M3 generally enhances existing model merging methods, achieving over 10-point improvements on multiple datasets and up to 30 points on one. For example, the improvements achieved by Average Merging with M3 for Math & Code are 7.43% on GSM8K, 3.74% on Math, and 11.0% on MBPP. For LM & Code, Average Merging with M3 shows improvements of 7.31% on AlpacaEval, 7.32% on HumanEval, and 2.4% on MBPP. Task Arithmetic with M3 results in improvements of 2.0% on AlpacaEval and 2.44% on HumanEval for LM & Code, and 10.4% on MBPP for Math & Code. TIES-Merging with M3 achieves an improvement of 4.01% for LM & Math on GSM8K. For LM & Code, TIES-Merging with M3 shows significant improvements of 3.11% on AlpacaEval, 25.61% on HumanEval, and 30.8% on MBPP. Furthermore, as evidenced by the Avg. column in Table 2, M3 consistently improves the overall performance of the merged models. This enhancement is particularly crucial for scenarios where the merged models must strike a balance between task-specific capabilities. A few datasets do not show performance improvements with M3, but the results remain on par with those without M3, indicating that it does not degrade overall performance. Moreover, such cases are negligible compared to the overall benefits brought by M3.
As shown in Table 2, M3 achieves the best performance on four out of five task-specific datasets, surpassing even the individually fine-tuned LLMs (i.e., LM, Math, and Code). This suggests that M3 enables the discovery of interpolated models along linear paths between two models that outperform both endpoints, providing support for the theoretical insights discussed in Section 3.2.
In addition, we also compare our proposed method M3 with existing model merging methods. The details are in Appendix G.
Mitigating Suboptimal Fine-Tuning with M3
Yu et al. (2024) noted that the poor performance of merging WizardMath-13B with llama-2-13b-code-alpaca stems from the latter not being sufficiently fine-tuned for code generation tasks. As shown in Table 2, the proposed M3 method improves the pass@1 score on the code generation dataset MBPP by 10.4%. This improvement demonstrates that even when one of the fine-tuned models is suboptimal for a specific task, M3 can effectively unlock the merging potential by adaptively exploring suitable interpolation coefficients (i.e., contribution ratios), thereby maximizing the performance of the merged model. In other words, M3 significantly mitigates the negative impact of insufficient fine-tuning on merging quality, further highlighting the critical importance of task-specific LLMs’ contribution ratio in determining the performance of the merged model.
@@ 4.3 Model Robustness
To ensure that the evaluation datasets are as representative as possible of OOD data, we select datasets with sufficiently recent release dates and ensure they cover domains that fine-tuned LLMs have not been specifically trained on. Consequently, Math & Code is evaluated on LiveBench-Instruction, LM & Math on LiveBench-Coding, and LM & Code on LiveBench-TypoFixing. As illustrated in Figure 4, M3 consistently enhances the performance of merged models—Math & Code, LM & Math, and LM & Code—on OOD datasets. Specifically, when Task Arithmetic is combined with M3, the Math & Code model demonstrates a 1.9% improvement in win rate on LiveBench-Instruction, the LM & Math model achieves a 1.6% increase in pass@1 on LiveBench-Coding, and the LM & Code model shows a significant 6% boost in accuracy on LiveBench-TypoFixing. Similarly, when Average Merging is combined with M3, the Math & Code model attains a 1.5% improvement in win rate on LiveBench-Instruction, the LM & Math model achieves a 0.7% increase in pass@1 on LiveBench-Coding, and the LM & Code model exhibits a 4% enhancement in accuracy on LiveBench-TypoFixing. Finally, when TIES-Merging is applied alongside M3, the Math & Code model achieves a 1.1% improvement in win rate on LiveBench-Instruction, the LM & Math model records a 0.6% increase in pass@1 on LiveBench-Coding, and the LM & Code model demonstrates a remarkable 14% improvement in accuracy on LiveBench-TypoFixing.
@@
We employ three Prompt Attack Methods supported by the promptbench codebase (DeepWordBug, BERTAttack, and StressTest) (Zhu et al., 2024) to evaluate the adversarial robustness of three merged models (Math & Code, LM & Math, and LM & Code) obtained through the task arithmetic method. To balance experimental effectiveness with computational efficiency, we randomly selected the positions of the three attacked words in the prompts when executing the DeepWordBug and BERTAttack attacks. Adversarial robustness is assessed using the Performance Drop Rate (PDR) (Zhu et al., 2023), where a lower PDR indicates stronger robustness. Further details on PDR can be found in Appendix D.
@@
4.4 Comparison of Mixup Model Merge and DARE
DARE is a model sparsification method proposed by (Yu et al., 2024), with a more detailed introduction provided in Appendix E. This section takes DARE as an example to discuss the relationship between the existing model sparsification method and M3. We integrated M3 and DARE with three existing model merging techniques—Average Merging, Task Arithmetic, and TIES-Merging—to compare these two plug-and-play methods. Detailed experimental results are provided in Appendix F. Additionally, in the DARE method, the drop rate hyperparameter is set to 0.2.
From Table 8, it can be concluded that M3 generally outperforms DARE, exhibiting significantly superior performance on certain datasets. For instance, the Math & Code model achieves a pass@1 score of 9.8% on the MBPP dataset when combined with DARE, while this score increases to 19% when combined with M3.
As shown in Table 8, Combining DARE and M3 generally results in better model merging performance. This suggests that M3 and DARE can complement each other. Moreover, in some cases, M3 alone can deliver the best results, while DARE alone only achieves the best performance in very few cases, further demonstrating the superiority of M3.
@@ 5 Conclusion
Addressing the issue that existing methods overlook the varying contribution ratios of different models, we propose Mixup Model Merge (M3), a novel approach based on the randomized linear interpolation strategy of Mixup. M3 performs linear interpolation in the parameter space between two task-specific large language models, with interpolation coefficients sampled from a Beta distribution to explore a diverse range of contribution ratios. This method leverages controllable randomness to significantly outperform traditional equal-ratio merging. Extensive experiments demonstrate that M3 not only improves merged model performance across multiple tasks but also enhances out-of-distribution and adversarial robustness, and effectively combines with model sparsification methods such as DARE. By tuning the Beta distribution parameters, M3 balances efficiency and diversity in exploring contribution ratios.
@@ 6 Limitations
Despite its advantages, the M3 method has several limitations: (1) Our method is currently limited to pairwise model merging, as random linear interpolation is typically defined between two entities. Extending it to merge three or more models requires new strategies for sampling multi-way interpolation coefficients, which introduces additional complexity and remains an open direction for future work. (2) This study focuses on merging homogeneous models fine-tuned from the same pretrained backbone. Applying M3 to heterogeneous models with different architectures or fine-tuning paths poses new challenges and will be explored in future work. (3) Our method may also be extended to broader applications, such as merging fine-tuned models with Reinforcement Learning with Human Feedback (RLHF) (Ouyang et al., 2022) models to reduce the alignment tax (FINE-TUNING, ).
@@ Appendix E Detailed Introduction to DARE
Appendix F Detailed M3 and DARE Comparison Results
Merging Methods Models Use M3 (Ours) Use DARE Instruction Following Mathematical Reasoning Code Generating
AlpacaEval GSM8K MATH HumanEval MBPP
Average Merging LM & Math Yes No 44.40 66.26 13.80 - -
No Yes 44.22 66.57 12.96 - -
Yes Yes 43.53 66.57 14.12 - -
LM & Code Yes No 43.91 - - 37.20 34.40
No Yes 38.81 - - 31.71 32.40
Yes Yes 40.31 - - 36.59 37.00
Math & Code Yes No - 63.61 14.02 8.54 19.20$\uparrow$
No Yes - 56.18 10.28 6.10 7.80
Yes Yes - 64.97 13.54 9.76 21.20
Task Arithmetic LM & Math Yes No 41.65 66.34 13.74 - -
No Yes 49.00 66.64 13.02 - -
Yes Yes 44.90 67.32 13.74 - -
LM & Code Yes No 46.64 - - 35.37 33.80
No Yes 41.47 - - 35.98 33.00
Yes Yes 45.20 - - 35.98 35.20
Math & Code Yes No - 63.53 13.94 7.93 19.00
No Yes - 65.05 13.96 10.37 9.80
Yes Yes - 65.13 14.32 8.54 18.00
TIES- Merging LM & Math Yes No 38.73 18.57 2.48 - -
No Yes 37.92 18.04 2.34 - -
Yes Yes 39.93 19.26 2.82 - -
LM & Code Yes No 44.96 - - 25.61 30.80
No Yes 43.13 - - 0.0 0.0
Yes Yes 45.65 - - 26.83 33.20
Math & Code Yes No - 64.75 14.16 9.76 21.4
No Yes - 64.82 13.88 10.37 23.60
Yes Yes - 64.75 14.78 9.15 19.60
Table 8: Comparison of our method M3 and DARE. The best and second-best results are marked in bold and underlined fonts.
We compare the two plug-and-play methods, M3 and DARE, by integrating each with three existing model merging techniques—Average Merging, Task Arithmetic, and TIES-Merging. The detailed experimental results are presented in Table 8.
Appendix G Comparison of M3 and Existing Model Merging Approaches
Models Merging Methods Instruction Following Mathematical Reasoning Code Generating Avg.
AlpacaEval GSM8K MBPP
LM & Math Average Merging 45.28 66.34 - 55.81
Task Arithmetic 45.78 66.34 - 56.06
TIES-Merging 38.63 14.56 - 26.60
DARE 49.00 66.64 - 57.82
DELLA 44.03 48.45 - 46.24
M3 (Ours) 44.90 67.32 - 56.11
LM & Code Average Merging 36.60 - 32.00 34.3
Task Arithmetic 44.64 - 33.60 39.12
TIES-Merging 41.85 - 0.0 20.93
DARE 41.47 - 33.00 37.24
DELLA 40.99 - 35.00 38.00
M3 (Ours) 45.20 - 35.20 40.2
Math & Code Average Merging - 56.17 8.20 32.19
Task Arithmetic - 64.67 8.60 36.64
TIES-Merging - 64.67 22.60 43.64
DARE - 65.05 9.80 37.43
DELLA - 63.08 20.2 41.64
M3 (Ours) - 64.75 21.4 43.10
Table 9: Comparison of our method M3 and existing model merging approaches. The best and second-best results are marked in bold and underlined fonts.
For each baseline method—Average Merging, Task Arithmetic, TIES-Merging, and DARE—we combine it with our proposed M3 method and select the best-performing combination for each merged model (LM & Math, LM & Code and Math & Code). These optimal results are reported as M3 (Ours) in Table 9.
As presented in Table 9, our proposed method M3 achieves the best overall performance across the three evaluated tasks. Importantly, M3 attains these improvements without involving complex architectural designs or intensive computations; it merely adjusts the contribution ratios of constituent models to realize substantial gains, thereby outperforming the DELLA approach. DELLA advances model sparsification techniques by building upon TIES-Merging and DARE, introducing a novel pruning strategy termed MAGPRUNE, which selects delta parameters based on their magnitudes. This strategy has inspired us to investigate new directions for enhancing the performance of merged models.